# Decomposition of Borel graphs and cohomology

### v2

@@ Abstract
1 Introduction
1.1 Main theorem
Accessibility. The Stallings theorem for ends of groups [Sta, Ber] states that a finitely generated group with more than one ends decomposes into either an amalgamated free product or an HNN extension over a finite subgroup. Dunwoody’s accessibility is the notion that requires this decomposition process to finish in finite steps.
@@ 1.1 Main theorem
Borel graphs. Let $X$ be a standard Borel space. A Borel equivalence relation on $X$ is an equivalence relation on $X$ which is a Borel subset of $X\times X$. It is called a countable Borel equivalence relation if each equivalence class is at most countable. They have been studied in the context of ergodic theory and descriptive set theory. A Borel graph on $X$ is a simplicial graph on $X$ whose edge set is a Borel subset of $X\times X$. If it is locally countable, then its connected relation defines a countable Borel equivalence relation on $X$. This is a generalization of the orbit equivalence relation associated with a Borel action of a countable group on $X$. One of the main theme of this area is to see how combinatorial or geometric structures of Borel graphs are reflected in the equivalence relations that they generate. Recently, there have been many attempts to apply ideas of geometric group theory to the study of Borel graphs (e.g., [CJM+],[CGMT]).
Tserunyan establishes an analog of the Stallings theorem for Borel graphs [Tse, Theorem 1.4] (see also [Ghy, Théorème D] and [Pau, Theorem C] for other analogs). It is proved that if a locally finite Borel graph $G$ has components with more than one ends, then the Borel equivalence relation generated by $G$ is a free product of two subequivalence relations, one of which is treeable, i.e., generated by a Borel acyclic graph.
@@
Here $R_{G}$ is an algebra associated to $G$, which is an analog of the group ring for a group. As explained below, it is defined only by the metric strucure of $(X,G)$ as an abstarct graph, that is, its measurable structure is forgotten. Hence the cohomological assumption of Theorem 1.3 is only about the metric structure. Under this assumption, the theorem states that $G$ is replaced by a Borel graph $G^{\prime}$, roughly preserving the metric strucure, and $G^{\prime}$ freely decomposes into Borel subgraphs $T$ and $H$ such that $T$ is acyclic and all components of $H$ are at most one-ended in a strong sense. In fact, the converse of this theorem also holds (Proposition 4.24).
@@
for every $10\leq i\leq n-$ 1, then we have $x_{0}\neq x_{2n}$.
@@
The algebra and cohomology. Let $R$ be a commutative ring.
@@
$G^{k}$ =$=\{(x,y)\in E_{G}\mid d_{G}(x,y)\leq k\}\textup{ and}$
$R_{G}^{k}$ =$=\left\{a\in l^{\infty}_{R}(E_{G})\>\middle|\>a\left(E_{G}\setminus G^{k}\right)=0\right\}.$
@@
where $[x]_{G}$ is the $E_{G}$-class containing $x$. This is well-defined since $G$ has uniformly bounded degrees.
@@
$G=\left\{\left(x,s^{\pm 1}x\right)\in X\times X\>\middle|\>x\in X,\ s\in S\right\}.$
@@
1.2 Borel graphs of cohomological dimension one
Recall that for a commutative ring $R$, the $R$-cohomological dimension of a group $\Gamma$, which is denoted by $\textup{cd}_{R}(\Gamma)$, is the smallest $n\in\{0,1,2,...,\infty\}$ such that $\textup{H}^{i}(\Gamma,M)=0$ for all left $R\Gamma$-module $M$ and $i>n$. One of the applications of the Stallings theorem for ends of groups is the Stallings-Swan theorem that for any non-zero commutative ring $R$, a torsion free group $\Gamma$ with $\textup{cd}_{R}(\Gamma)\leq 1$ must be free [Swa, Theorem A]. By using accessibility of groups, Dunwoody gives a variant of this theorem as follows:
Theorem 1.12 ([Dun, Corollary 1.2]).
@@ 1.2 Borel graphs of cohomological dimension one
Let $R$ be a commutative ring and $(X,G)$ a graph with uniformly bounded degrees. The $R$-cohomological dimension of $G$, which is denoted by $\textup{cd}_{R}(G)$, is the smallest $n\in\{0,1,2,...,\infty\}$ such that $\textup{H}^{i}(G,M)=0$ for all left $R_{G}$-module $M$ and $i>n$.
Then as a consequence of Theorem 1.3, we obtain an analog of Theorem 1.12 for Borel graphs as follows:
Theorem 1.15.
@@
Here, we say that a graph $H$ on $X$ is Lipschitz equivalent to $G$ if $\textup{id}_{X}:(X,H)\to(X,G)$ is a quasi-isometry.
An application. Treeable equivalence relations have been studied as an analog of free groups. Chen-Poulin-Tao-Tserunyan show that treeability has a property similar to the quasi-isometric rigidity of free groups: If $(X,G)$ is a locally finite Borel graph such that every component is quasi-isometric to a tree, then the equivalence relation generated by $G$ is treeable [CPTT, Theorem 1.1]. Their proof applies the theory of median graphs to Borel graphs. Moreover, by refining the argument, they show the following:
Theorem 1.16 ([CPTT, Theorem 1.2]).
@@
Now we give an alternate proof of this theorem through Theorem 1.15: Let $R$ be a non-zero commutative ring. By assumption, there exists a (not necessarily Borel) acyclic graph $T$ on $X$ Lipschitz equivalent to $G$. Note that $R_{G}=R_{T}$ holds. Since $T$ is acyclic, it has $R$-cohomological dimension at most $1$ (Lemma 4.9), and so does $G$. Then by Theorem 1.15, there exists a Borel acyclic graph on $X$ Lipschitz equivalent to $G$. Theorem 1.16 is proved.
This new proof has a common feature with the standard proof of the quasi-isometric rigidity of free groups. Indeed, our proof relies on the decomposition of Borel graphs given by Theorem 1.3 in the same way that the proof given in [GD, Chapitre 7, Théorème 19] (see also [DK, Theorem 20.45]) relies on accessibility of groups.
@@
In [CPTT, Theorem 1.4], it is also proved that if $(X,G)$ is a locally finite Borel graph such that every component has bounded tree-width, then the equivalence relation generated by $G$ is treeable. Jardón-Sánchez also proves this result in the case of uniformly bounded degrees, by using tree decompositions [Jar, Theorem 3]. His method is similar to ours in some ways, and can be applied to the setting of Theorem 1.16 since a bounded-degree quasi-tree has bounded tree-width. However it is non-trivial to make the resulting Borel acyclic graph Lipschitz equivalent to $G$. See also Remark 5.4.
Organization. In section 2, we prepare basic terminology and facts on graph theory and homological algebra. In section 3, we present a construction of structure trees, and explain how to use it for Borel graphs. In section 4, we investigate the cohomology of graphs, in particular, the group $\textup{H}^{1}(G,R_{G})$. Then Theorems 1.3 and 1.15 are proved in section 5.
2 Preliminaries
2.1 Terminology of graph theory
@@
  • If a sequence $\{x_{i}\}_{i=0}^{n}$ satisfies $(x_{i},x_{i+1})\in G$ for every $10\leq i\leq n-$ 1, then it is called a ($G$-)path from $x_{0}$ to $x_{n}$. The length of this path is $n$. This path is simple if $x_{i}\neq x_{j}$ for all $i\neq j$, and is a cycle if $x_{0}=x_{n}$.
@@
  • The degree of $x\in X$ is the number |$|\{y\in X\mid(x,y)\in G\}|$. The graph $(X,G)$ has uniformly bounded degrees if $\sup_{x\in X}|\{y\in X\mid(x,y)\in G\}|<\infty$ holds.
@@
    For $x\in X$, the $E_{G}$-class containing $x$ is denoted by $[x]_{G}$. For $k\geq 0$, set
@@
  • Let $G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}⟨$⟩ be the set of bijections $\varphi:\textup{dom}\varphi\to\textup{im}\varphi$ between subsets of $X$ such that $\sup_{x\in\textup{dom}\varphi}d_{G}(\varphi x,x)<\infty$. The composition and the inverse on $G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}⟨$⟩ are naturally defined. For $\varphi\in\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩, set
@@
Let $(X,G)$ be a graph with uniformly bounded degrees. Then there exists a family $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\leavevmode\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\leavevmode\hbox{${\rangle}$}}$⟩ such that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$.
@@
$\partial_{\textup{iv}}^{G}C$ =$=\left\{x\in C\>\middle|\>\exists y\in\overline{C},\ (x,y)\in G\right\},$
@@
$\partial_{\textup{ie}}^{G}C$ $=\left(\overline{C}\times C\right)\cap G\textup{ and}$
@@
2.2 Basic homological algebra
@@
An $R$-projective resolution of a left $R$-module $N$ is an exact sequence
[…]
@@
[…]
@@
The projective dimension of a left $R$-module $N$ is the smallest $n\in\{0,1,2,...,\infty\}$ satisfying any of conditions (i)-(iv). It is denoted by $\textup{pd}_{R}(N)$.
@@
A left $R$-module $N$ is projective if and only if $\textup{pd}_{R}(N)=0$.
@@
be an $R$-projective resolution. Since $\textup{Ext}_{R}^{n}(N,R)=0$ holds and the functor $\textup{Ext}_{R}^{n}(N,-)$ preserves direct sums, we have $\textup{Ext}_{R}^{n}(N,F)=0$. This implies that the map $-\circ\partial_{n}:\textup{Hom}_{R}(P_{n-1},F)\to\textup{Hom}_{R}(P_{n},F)$ is surjective. Also the map $q\circ-:\textup{Hom}_{R}(P_{n},F)\to\textup{Hom}_{R}(P_{n},M)$ is surjective since $P_{n}$ is projective. Now the following diagram commutes
@@
Then since the upper $-\circ\partial_{n}$ and the right-side $q\circ-$ are surjective, the lower $-\circ\partial_{n}$ is also surjective. Hence we have $\textup{Ext}_{R}^{n}(N,M)=0$ for every left $R_{G}$-module $M$. This means that $\textup{pd}_{R}(N)\leq n-1$. ∎
@@
Suppose that a left $R$-module $N$ is projective and
[…]
@@
for $n\geq 1$, and these are 00 since $N$ is projective. The exactness at $M\otimes_{R}P_{0}$ and $M\otimes_{R}N$ is trivial since the functor $M\otimes_{R}-$ is right exact. ∎
3 Structure trees
3.1 Construction of trees

### v3

@@ Abstract.
1. Introduction
1.1. Decomposition of Borel graphs
Accessibility of groups
The Stallings theorem for ends of groups [Sta68, Ber68] states that a finitely generated group with more than one ends decomposes into either an amalgamated free product or an HNN extension over a finite subgroup. Dunwoody’s accessibility is the notion that requires this decomposition process to finish in finite steps.
@@ 1.1. Decomposition of Borel graphs
Borel graphs.
Let $X$ be a standard Borel space. A Borel equivalence relation on $X$ is an equivalence relation on $X$ which is a Borel subset of $X\times X$. It is called a countable Borel equivalence relation if each equivalence class is at most countable. They have been studied in the context of ergodic theory and descriptive set theory. A Borel graph on $X$ is a simplicial graph on $X$ whose edge set is a Borel subset of $X\times X$. If it is locally countable, then its connected relation defines a countable Borel equivalence relation on $X$. This is a generalization of the orbit equivalence relation associated with a Borel action of a countable group on $X$. One of the main theme of this area is to see how combinatorial or geometric structures of Borel graphs are reflected in the equivalence relations that they generate. Recently, there have been many attempts to apply ideas of geometric group theory to the study of Borel graphs (e.g., [CJM+23], [CGM+21]).
Tserunyan establishes an analog of the Stallings theorem for Borel graphs [Tse20, Theorem 4.1] (see also [Ghy95, Théorème D] and [Pau99, Theorem C] for other analogs). It is proved that if a locally finite Borel graph $G$ has components with more than one ends, then the Borel equivalence relation generated by $G$ is a free product of two subequivalence relations, one of which is treeable, i.e., generated by a Borel acyclic graph.
@@
Here $R_{G}$ is an algebra associated to $G$, which is an analog of the group ring for a group. As explained below, it is defined only by the metric strucure of $(X,G)$ as an abstarct graph, that is, its measurable structure is forgotten. Hence the cohomological assumption of Theorem A is only involved with the metric structure.
The consequence of the theorem states that $G$ admits a nice decomposition into two Borel graphs $T$ and $H$, which is an analog of Bass-Serre decomposition of a group. Condition (i) and (ii) means that a vertex of the Bass-Serre tree correspond to an $E_{H}$-class and an edge of the Bass-Serre tree correspond to an edge of $T$. Finally, condition (iii) means that the decomposition is optimal, i.e., $H$ cannot be decomposed essentially anymore.
There are two major contributions of us. First, we not only give a free product decomposition $E_{1}\ast E_{2}$, but also give two Borel graphs $T$ and $H$ generating $E_{1}$ and $E_{2}$ respectively such that the union of $T$ and $H$ is “close” to $G$. To be precise, the construction of $T$ basically follows [Tse20, Theorem 4.1], but as far as we understand, the construction $H$ appears first in this paper (see Proposition 3.17). Second, we establish the notion of “optimal decomposition” and characterize when it is possible. This is a crucial part of the paper, where cohomology theory plays a role. Further, the optimality of the decomposition leads to an application in section 1.2.
Remark 1.3 (Proposition 4.24).
The converse of Theorem A also holds. This is not in the context of Borel graph theory, but assures that the assumption of Theorem A is proper.
@@
for every $0\leq i\leq n-1$, then we have $x_{0}\neq x_{2n}$.
@@
The algebra and cohomology.
Let $R$ be a commutative ring.
@@
$G^{k}$ $=\{(x,y)\in E_{G}\mid d_{G}(x,y)\leq k\}\textup{ and}$
$R_{G}^{k}$ $=\left\{a\in l^{\infty}_{R}(E_{G})\>\middle|\>a(E_{G}\setminus G^{k})=0\right\}.$
@@
where $x]_{G}$ is the $E_{G}$-class containing $x$. This is well-defined since $G$ has uniformly bounded degrees.
@@
$G=\left\{(x,s^{\pm 1}x)\in X\times X\>\middle|\>x\in X,\ s\in S\right\}.$
@@
Remark 1.12.
We explain the relationship with Roe’s coarse cohomology [Roe03, Chapter 5]. Note that in the above setting, the path metric $d_{G}$ defines the canonical coarse structure on $X$, so the coarse cohomology $\textup{HX}^{n}(X,R)$ can be defined. Then we have the isomorphism
$\textup{HX}^{n}(X,R)\simeq\bigoplus_{x\in A}\textup{H}^{n}(G,R_{G})\cdot 1_{\{x\}},$
where $A\subset X$ is any representative set for the equivalence relation $E_{G}$. These cohomology groups have similar properties, but $\textup{H}^{n}(G,R_{G})$ is involved with more uniform conditions than $\textup{HX}^{n}(X,R)$ in general. For example, $\textup{HX}^{1}(X,R)=0$ if and only if every component of $G$ is at most one-ended. As a result, if $(X,G)$ is a one-ended tree, then $\textup{HX}^{1}(X,R)=0$ but $\textup{H}^{1}(G,R_{G})\neq 0$. Another important difference is that $\textup{H}^{n}(G,R_{G})$ has a structure of right $R_{G}$-module while $\textup{HX}^{n}(X,R)$ does not admit such an action of translations. This enables us to discuss the “finite generation” condition for $\textup{H}^{1}(G,R_{G})$ in an appropriate sense.
1.2. Borel graphs of cohomological dimension one
Recall that for a commutative ring $R$, the $R$-cohomological dimension of a group $\Gamma$, which is denoted by $\textup{cd}_{R}(\Gamma)$, is the smallest $n\in\{0,1,2,$$\infty\}$ such that $\textup{H}^{i}(\Gamma,M)=0$ for all left $R\Gamma$-module $M$ and $i>n$. One of the applications of the Stallings theorem for ends of groups is Stallings-Swan’s theorem that for any non-zero commutative ring $R$, a torsion free group $\Gamma$ with $\textup{cd}_{R}(\Gamma)\leq 1$ must be free [Swa69, Theorem A]. By using accessibility of groups, Dunwoody gives a variant of this theorem as follows:
Theorem 1.13 ([Dun79, Corollary 1.2]).
@@ 1.2. Borel graphs of cohomological dimension one
Let $R$ be a commutative ring and $(X,G)$ a graph with uniformly bounded degrees. The $R$-cohomological dimension of $G$, which is denoted by $\textup{cd}_{R}(G)$, is the smallest $n\in\{0,1,2,$$\infty\}$ such that $\textup{H}^{i}(G,M)=0$ for all left $R_{G}$-module $M$ and $i>n$.
Then as a consequence of Theorem A, we obtain an analog of Theorem 1.13 for Borel graphs as follows:
Theorem B (Theorem 5.5).
@@
Here, we say that a graph $H$ on $X$ is Lipschitz equivalent to $G$ if $\textup{id}_{X}:(X,H)\to(X,G)$ is a quasi-isometry. We should emphasize again that condition (i) ensures the existence of another Borel graph even though condition (ii) is only concerning the metric structure.
An application.
Treeable equivalence relations have been studied as an analog of free groups. Chen-Poulin-Tao-Tserunyan show that treeability has a property similar to the quasi-isometric rigidity of free groups: If $(X,G)$ is a locally finite Borel graph such that every component is quasi-isometric to a tree, then the equivalence relation generated by $G$ is treeable [CPT+25, Theorem 1.1]. Their proof applies the theory of median graphs to Borel graphs. Moreover, by refining the argument, they show the following:
Theorem 1.16 ([CPT+25, Theorem 1.2]).
@@
Now we give an alternate proof of this theorem through Theorem B: Let $R$ be a non-zero commutative ring. By assumption, there exists a (not necessarily Borel) acyclic graph $T$ on $X$ Lipschitz equivalent to $G$. Note that $R_{G}=R_{T}$ holds. Since $T$ is acyclic, it has $R$-cohomological dimension at most $1$ (Lemma 4.9), and so does $G$. Then by Theorem B, there exists a Borel acyclic graph on $X$ Lipschitz equivalent to $G$. Theorem 1.16 is proved.
This new proof has a common feature with the standard proof of the quasi-isometric rigidity of free groups. Indeed, our proof relies on the decomposition of Borel graphs given by Theorem A in the same way that the proofs of [GH90, Chapitre 7, Théorème 19] and [DK18, Theorem 20.45] rely on accessibility of groups.
@@
[…]
Remark 1.18.
It is proved in [CPT+25, Theorem 1.4] that if $(X,G)$ is a locally finite Borel graph such that every component has bounded tree-width, then the equivalence relation generated by $G$ is treeable. Jardón-Sánchez also give a proof of this result in the case of uniformly bounded degrees, by using tree decompositions [Jar23, Theorem 3]. This method has some similarity to ours, and can be applied to the setting of Theorem 1.16 since a bounded-degree quasi-tree has bounded tree-width. However, it is non-trivial to make the resulting Borel acyclic graph Lipschitz equivalent to $G$. See also Remark 5.4.
1.3. Outline of the paper
In section 2, we prepare basic terminology and facts on graph theory and homological algebra.
In section 3, we present a construction of the structure tree associated to a treeset, which is a main tool to prove our main theorem. Then we apply it to Borel graphs and obtain a prototype of the decomposition.
In section 4, we first describe properties of the cohomology and the cohomological dimension of graphs. Then we investigate the group $\textup{H}^{1}(G,R_{G})$, and its relation with the cutsets of $G$.
In section 5, by combining all the above content, we prove Theorems A and B.
2. Preliminaries
2.1. Terminology of graph theory
@@
  • If a sequence $\{x_{i}\}_{i=0}^{n}$ satisfies $(x_{i},x_{i+1})\in G$ for every $0\leq i\leq n-1$, then it is called a ($G$-)path from $x_{0}$ to $x_{n}$. The length of this path is $n$. This path is simple if $x_{i}\neq x_{j}$ for all $i\neq j$, and is a cycle if $x_{0}=x_{n}$.
@@
  • The degree of $x\in X$ is the number $|\{y\in X\mid(x,y)\in G\}|$. The graph $(X,G)$ has uniformly bounded degrees if $\sup_{x\in X}|\{y\in X\mid(x,y)\in G\}|<\infty$ holds.
@@
    For $x\in X$, the $E_{G}$-class containing $x$ is denoted by $x]_{G}$. For $k\geq 0$, set
@@
[…]
@@
Let $(X,G)$ be a graph with uniformly bounded degrees. Then there exists a family $\{\varphi_{i}\}_{i=1}^{n}\subset\mathopen{\hbox{${\langle}$}\kern-1.94444pt\hbox{${\langle}$}}G\mathclose{\hbox{${\rangle}$}\kern-1.94444pt\hbox{${\rangle}$}}$ such that $G=\bigsqcup_{i=1}^{n}\left(\textup{graph}\varphi_{i}\sqcup\textup{graph}\varphi_{i}^{-1}\right)$.
@@
$\partial_{\textup{iv}}^{G}C$ $=\left\{x\in C\>\middle|\>\exists y\in\overline{C},\ (x,y)\in G\right\},$
@@
$\partial_{\textup{ie}}^{G}C$ $=(\overline{C}\times C)\cap G\textup{ and}$
@@
2.2. Basic homological algebra
@@
An $R$-projective resolution of a left $R$-module NN is an exact sequence
$\cdots\rightarrow P_{n}\xrightarrow{\partial_{n}}P_{n-1}\to\cdots\xrightarrow{\partial_{1}}P_{0}\xrightarrow{\varepsilon}N\rightarrow 0$
@@
$\cdots\rightarrow P_{n}\xrightarrow{\partial_{n}}P_{n-1}\to\cdots\xrightarrow{\partial_{1}}P_{0}\xrightarrow{\varepsilon}N\rightarrow 0.$
@@
The projective dimension of a left $R$-module NN is the smallest $n\in\{0,1,2,$$\infty\}$ satisfying any of conditions (i)-(iv). It is denoted by $\textup{pd}_{R}(N)$.
@@
A left $R$-module NN is projective if and only if $\textup{pd}_{R}(N)=0$.
@@
be an $R$-projective resolution. Since $\textup{Ext}_{R}^{n}(N,R)=0$ holds and the functor $\textup{Ext}_{R}^{n}(N,-)$ preserves direct sums, we have $\textup{Ext}_{R}^{n}(N,F)=0$. This implies that the map $\circ\partial_{n}:\textup{Hom}_{R}(P_{n-1},F)\to\textup{Hom}_{R}(P_{n},F)$ is surjective. Also the map $q\circ-:\textup{Hom}_{R}(P_{n},F)\to\textup{Hom}_{R}(P_{n},M)$ is surjective since $P_{n}$ is projective. Now the following diagram commutes
@@
Then since the upper $\circ\partial_{n}$ and the right-side $q\circ$- are surjective, the lower $\circ\partial_{n}$ is also surjective. Hence we have $\textup{Ext}_{R}^{n}(N,M)=0$ for every left $R_{G}$-module $M$. This means that $\textup{pd}_{R}(N)\leq n-1$. ∎
@@
Suppose that a left $R$-module NN is projective and
$\cdots\to P_{n}\xrightarrow{\partial_{n}}P_{n-1}\to\cdots\to P_{0}\xrightarrow{\varepsilon}N\to 0$
@@
for $n\geq 1$, and these are 0 since $N$ is projective. The exactness at $M\otimes_{R}P_{0}$ and $M\otimes_{R}N$ is trivial since the functor $M\otimes_{R}$- is right exact. ∎
3. Structure trees
The goal of this section is to show Proposition 3.17.
3.1. Construction of trees